# Optimal and Fair Encouragement Policy Evaluation and Learning

### v4

@@
Abstract
In consequential domains, it is often impossible to compel individuals to take treatment, so that optimal policy rules are merely suggestions in the presence of human non-adherence to treatment recommendations. Under heterogeneity, covariates may predict take-up of treatment and final outcome, but differently. While optimal treatment rules optimize causal outcomes across the population, access parity constraints or other fairness considerations on who receives treatment can be important. For example, in social services, a persistent puzzle is the gap in take-up of beneficial services among those who may benefit from them the most. We study causal identification and robust estimation of optimal treatment rules, including under potential violations of positivity. We consider fairness constraints such as demographic parity in treatment take-up, and other constraints, via constrained optimization. Our framework can be extended to handle algorithmic recommendations under an often-reasonable covariate-conditional exclusion restriction, using our robustness checks for lack of positivity in the recommendation. We develop a two-stage algorithm for solving over parametrized policy classes under general constraints to obtain variance-sensitive regret bounds. We illustrate the methods in three case studies based on data from reminders of SNAP benefits recertification, randomized encouragement to enroll in insurance, and from pretrial supervised release with electronic monitoring. While the specific remedy to inequities in algorithmic allocation is context-specific, it requires studying both take-up of decisions and downstream outcomes of them.
@@ 1 Introduction
Combining causal inference and machine learning to estimate heterogeneous treatment effects can increase revenue and improve public and individual health outcomes by facilitating personalized treatments such as medication decisions or interactions with e-commerce platforms and targeting of social interventions to those who benefit from them most. In particular, a large literature on so-called optimal treatment regimes, or off-policy learning, studies how to leverage treatment effect heterogeneity to identify optimal personalized decision rules for prescriptive decisions about who ought to receive which treatment (Athey, 2017; Kitagawa and Tetenov, 2015; Manski, 2005; Zhao et al., 2012).
@@
However, in many important settings, we cannot compel individuals into treatment. A so-called encouragement design is used to randomize over encouragements, or recommendations into treatment, since it is impossible to randomize the treatment itself. Randomizing treatment may be not possible either because it is ultimately a human-in-the-loop who receives the treatment and it is impossible to compel someone into treatment or because it would be unethical to withhold access to the treatment intervention from the control group. For example, in e-commerce, although companies can assign different users to different visual interfaces of a website, they cannot compel users to sign up for certain services: Instead, they can nudge or encourage users to sign up via promotions and offers. Alternatively, companies may not want to withhold access to a new feature to construct a control group and may instead randomize a new feature’s visibility or the extent of advertising for it (Spotify Engineering, 2023). Then, the treatment group receives more encouragement to use the intervention, while the control group has access but does not receive additional encouragement. Encouragement designs are pervasive not only in e-commerce but more generally in impact assessments across social settings, such as healthcare and social policy. Guidelines for the Abdul Latif Jameel Poverty Action Lab (J-PAL), a major provider of randomized controlled trials in development, note that “when evaluating a noncompulsory, entitlement program, researchers and practitioners cannot and should not force individuals to take up the program nor deny eligible individuals access to the program" and recommend encouragement designs (Heard et al., 2017).
The wide prevalence of encouragement designs implies that optimal treatment regimes also typically operate in the space of encouragements rather than treatment interventions themselves. In particular, we focus on the implications of personalized encouragement interventions for decision-making and fairness or equity concerns; we term such prospective decision rules optimal encouragement policies. We discuss methodological connections with and differences from the well-studied problem of nonadherence in causal inference later on. Often, for the purposes of decision-making, typical approaches adopt an intention-to-treat (ITT) perspective and take the encouragement itself as the treatment (i.e., directly optimizing downstream outcomes of encouragements). For example, pure pricing and profit maximization maximize the ITT effect (i.e., expected profit).
Optimal encouragement policies that personalize encouragement based on covariates are algorithmic allocations that ultimately induce treatment and outcomes. Since we learn about these allocations from historical data, they may suffer issues of algorithmic bias (Barocas et al., 2018). We generally take a multiobjective approach and consider tradeoffs between efficacy (outcomes induced by an optimal encouragement policy) and other equity measures, such as disparities in access (treatment takeup) or encouragement rates across protected groups. ITT-based analysis hides these multiobjective tradeoffs. For example, because of a long history of civil rights oversight, equitable access is of particular concern in public-sector allocations.
Methodologically, the key identifying assumption in evaluations of encouragement designs is what we call the covariate-conditional exclusion restriction (informally, the requirement that the encouragement impact outcomes through no channel other than by increasing treatment takeup), which specifically characterizes this type of design in our view. We give three motivating cross-sector examples that highlight the design’s distinguishing features, namely: 1) Firms can in practice act only in the space of encouragements, not that of treatments themselves; 2) encouragements affect outcomes only by modifying treatment probabilities; 3) firms might have fairness/equity concerns that require multiobjective comparisons; and 4) firms have the operational capacity to potentially target/personalize future encouragements under realistic resource constraints.
Example 1 (Pricing: Demand vs. revenue vs. long-term outcomes).
Firms may consider price discounts for the purposes of long-term outcomes via lead generation or customer acquisition. Price discounts could impact not only short-term revenue but also overall engagement and customer lifetime value. (Karlan and Zinman, 2008) study the impact of price discounts offered by a large microcredit lender. They randomize prices and find that lower interest rates would marginally decrease profits, but they also find evidence of heterogeneity: Women and lower-income individuals take up loans more under discounted rates, with no increase in default. For microcredit lenders interested in financial inclusion, personalized price discounts could expand access while minimizing profit reductions.
Example 2 (Healthcare: Adherence vs. treatment efficacy).
Nonadherence is a major issue in medication treatments and can be significantly affected by social determinants of health.
For example, Miao et al. (2024) studies reasons why patients might switch contraceptives (i.e., move off an initial prescription) and finds significantly higher switching rates for Black and Latinx groups because of not only medical concerns about side effects but also insurance coverage.
Our next example and case study introduces public-sector allocation questions regarding administrative burden in operations, which we revisit in our case study, so we provide additional context and go into more detail. The term administrative burden refers to costs associated with applying for, receiving, and participating in government benefits and services (OMB, 2022). There has been great policy and public-sector interest in reducing administrative burdens. These operational frictions can disproportionately disenfranchise marginalized and disadvantaged groups and can undermine equity-related policy goals. We model several different strategies to reduce administrative burden—reductions in information costs, improvements to outreach via omnichannel and digital marketing, and digital service design (OMB, 2022)—generally through encouragement into treatment (takeup of services).
Example 3 (Takeup of social services and digital outreach).
Many who are eligible for the Supplemental Nutrition Assistance Program (SNAP, formerly known as food stamps) do not receive benefits because of cumbersome processes and understaffed processing centers. While early economic models (Nichols and Zeckhauser, 1982) suggest that those who forgo SNAP application do so because they face a higher time opportunity cost of complying with administrative hassle and thus are likely to be ineligible or to qualify for only small amounts, empirically, this is not the case: Many who are denied benefits because of operational frictions (missed interviews, application errors) are actually eligible. “In 2019, one third of all applications in Los Angeles County were denied due to a missed interview," while five times as many were denied for ineligibility. (Finkelstein and Notowidigdo, 2019) examine behavioral frictions in SNAP benefit access procedures, randomizing both information about likely eligibility and higher-touch outreach via access to a counselor. Though they find overall positive average treatment effects, they also find that "the individuals who apply … in response to either intervention receive lower benefits and are less sick than the average enrollee in the control group”—that is, the applicants for whom the treatment had the intended effect were better off to begin with. How can we instead ensure that interventions reach the worst off, who might benefit most? Under resource constraints, tailoring who receives basic vs. more extensive and expensive outreach can help balance equity with efficiency. Koenecke et al. (2023) find that an efficient advertising budget for GetCalfresh, which streamlines enrollment in California’s CalFresh (SNAP), resulted in low takeup among Hispanic individuals: The authors conduct a survey and find general public support for equitable advertising budget allocations with higher spending on advertising aimed at harder-to-reach populations.
Later, we revisit data from Homonoff and Somerville’s (2021) study of text message reminders about recertification, i.e., the annual benefit renewal procedure. By leveraging heterogeneous effect estimation, we find that the intervention’s efficiency impact is broadly aligned with equity objectives, in the sense that the recipients for whom the treatment effects are greater in terms of next-year benefit dollars are positively rank-correlated with those more likely to respond to the text message reminders. However, personalization should be implemented with caution since this correlation is weaker for nonwhite than for white recipients, giving rise to large access gaps at the top of the distribution.
@@
We develop general methodology that models this structure under a set of causal assumptions, described in full detail later on. Thus far, we have focused on modeling physically randomized encouragements such as those we have just described. However, our model can be extended to model the impacts of algorithmic advice on humans-in-the-loop. In this algorithmic advice setting, the covariate-conditional exclusion restriction is still plausible, but the “overlap" assumption is violated: Randomized (or as-if randomized) algorithmic encouragement or nudge recommendations are typically fixed functions of covariates.
Example 4 (Algorithmic advice).
@@
Doctors prescribe treatment on the basis of algorithmic recommendations (Lin et al., 2021), managers and workers combine their expertise to act on the basis of algorithmic decision support (Bastani et al., 2021), and in the social sector, caseworkers assign individuals to benefit programs on the basis of recommendations from risk scores that support triage (De-Arteaga et al., 2020; Green and Chen, 2019; Yacoby et al., 2022).
Crucially, our model specifically disentangles the effects of algorithmic recommendations on final treatment decisions from the effects of the treatment on outcomes, a contribution that can help make recent approaches based on robustness less conservative. If, as in the settings that we consider here, human decision-makers have wide discretion resulting in treatment overlap, then rather than extrapolating the entire treatment effect function itself, we need extrapolate only the treatment response to the recommendation, which can be more plausible.
Our contributions are as follows. We characterize optimal and resource fairness–constrained optimal decision rules and develop statistically improved estimators and robustness checks for the setting of algorithmic recommendations with sufficiently randomized decisions. In contrast, previous work on algorithmic accountability focuses primarily on auditing recommendations rather than on both the access and efficacy achieved under the final decision rule. Therefore, previous methods can fall short in mitigating potential disparities. We consider two settings: one related to encouragement designs with random allocation of encouragement and another related to algorithmic recommendations (which requires either parametric or robust extrapolation). We also develop methodology for optimizing over a constrained policy class with less conservative out-of-sample fairness constraint satisfaction by means of a two-stage procedure, and we provide sample complexity bounds. We assess the improved recommendation rules in a stylized case study of optimization of health insurance expansion, using data from the Oregon Insurance study, and another stylized case study of optimization of recommendations for pretrial supervised release on the basis of a risk-assessment tool while reducing surveillance disparities.
@@ 2 Related Work
Optimal encouragement designs/policy learning with constraints. There is extensive literature on off-policy evaluation and learning, empirical welfare maximization, and optimal treatment regimes (Athey and Wager, 2021; Zhao et al., 2012; Manski, 2005; Kitagawa and Tetenov, 2015). Qiu et al. (2021) study an optimal individualized encouragement design, though their focus is on optimal individualized treatment regimes with instrumental variables (IVs). Kallus and Zhou (2021a) study fairness in pricing, and some of the desiderata in that setting on revenue (here, marginal welfare) and demand (takeup) are again relevant in our context, but in a more general setting beyond pricing. The results in Kallus and Zhou (2021a) are restricted to pricing and revenue maximization. We consider broader objectives. The most closely related work in terms of problem setup is the formulation of “optimal encouragement designs" in (Qiu et al., 2021). However, they focus on knapsack resource constraints, which have a different solution structure than fairness constraints. Sun et al. (2021) has studied uniform feasibility in constrained resource allocation, but without encouragement or fairness. Ben-Michael et al. (2021) studies robust extrapolation in policy learning from algorithmic recommendation, but not fairness. Our later case study on supervised release leverages the considerable randomness in final treatment decisions for supervised release (decisions that are arguably less consequential than pretrial detention decisions and hence subject to wide discretion) so that we require robust extrapolation over only the first out of two stages.
Fairness constraints in ITT analyses. We focus on deriving estimators for ITT analyses with an eye to relevant fairness constraints. Our interest is in imposing separate desiderata on treatment realizations under noncompliance, but we do not conduct IV inference: We assume that unconfoundedness holds. Our analysis essentially considers simultaneously two perspectives in the constrained optimization: 1) viewing treatment as a potential outcome of a recommendation treatment, i.e., $T(R)$, and 2) taking an ITT stance on the causal effects of treatment on outcomes, i.e., $Y(T)$, even though treatment is not controllable. Marginally, the first perspective yields estimates of disparities and is relevant for estimating fairness constraints, while the second is relevant for estimating the utility objective. Importantly, the quantities that we estimate are not on joint events of takeup and final outcome utility (in contrast to principal stratification). Rather, we assess personalized policies by their population-averaged utility and fairness measures.
Algorithmic advice in operations
A large and rapidly growing literature in operations studies algorithmic advice, at times from a behavioral focus (but not specialized to causal effects). (Ge et al., 2025) studies classification conditional parity under no-information on compliance, and proposes compliance-robust policies sandwiched in between the range of human decisions (under no structure as to how human responders behave). (Grand-Clément and Pauphilet, 2024) also shares a focus on leveraging adherence to optimize advice rather than decisions - however they adopt a dynamic model with updating. We focus on implications for optimizing causal encouragements as well. (McLaughlin and Spiess, 2022; Gillis et al., 2021) study the fairness of machine-assisted human decisions and develop a principal-agent model studying how algorithmic advice might shift beliefs in risk and preferences. In contrast, we are not micro-founded as to why disparities in compliance arise, which is an important direction of future work. On the non-algorithmic advice, but encouragement side, (Freund and Hssaine, 2025) consider dynamic resource allocation under fairness constraints, when deciding monetary incentives for retention under stochastic participation. Other models of stochastic resource usage and demand are also relevant, though potentially more specialized than our non-adherence setting.
3 Problem Setup
We briefly describe the problem setup. We work in the Neyman-Rubin potential outcomes framework for causal inference (Rubin, 1980). We define the following:

### v5

# MNSC-0001-2024.00

@@
[…]
MNSC-0001-2024.00
[…]
Optimal and Equitable Encouragement Policies
[…]
Mind the Gap: Optimal and Equitable Encouragement Policies
[…]
In consequential domains, it is often impossible to compel individuals to take treatment, so that optimal policy rules are merely suggestions in the presence of human non-adherence to treatment recommendations. We study personalized decision problems in which the planner controls recommendations into treatment rather than treatment itself. Under a covariate-conditional no-direct-effect model of encouragement, policy value depends on two distinct objects: responsiveness to encouragement and treatment efficacy. This modeling distinction makes induced treatment take-up, rather than recommendation rates alone, the natural fairness target and yields tractable policy characterizations under budget and access constraints. In settings with deterministic algorithmic recommendations, the same model localizes overlap-robustness to the recommendation-response model rather than the downstream outcome model. We illustrate the methods in case studies based on data from reminders of SNAP benefits recertification, and from pretrial supervised release with electronic monitoring. While the specific remedy to inequities in algorithmic allocation is context-specific, it requires studying both take-up of decisions and downstream outcomes of them.
[…]
@@ 1 Introduction
Data-driven decisions leverage individual-level data to learn heterogeneous treatment effects and can improve upon uniform decision rules for the population. Personalized targeting of promotions to customers can improve profits and longer-term revenues beyond one-size-fits-all campaigns (Simester et al. 2020). A crucial operational lever comprises of nudges and reminders, (Hirano et al. 2000, Karlan et al. 2016), and recent work shows that targeting nudges can further improve outcomes under a lower budget (von Zahn et al. 2025). A rapidly growing research literature studies personalized decision rules based on treatment effect heterogeneity to prescribe who ought to receive different treatments, deployed in healthcare, e-commerce, and social and societal settings (Athey 2017, Kitagawa and Tetenov 2015, Manski 2005, Zhao et al. 2012).
@@
However, in many important settings, we cannot compel individuals into taking treatment. Such mismatches between treatment assignment and realization are called non-compliance or non-adherence in the causal literature. In fact, non-adherence is inherent in encouragement designs that randomize over encouragements, or recommendations into treatment, when it is impossible to randomize the treatment itself due to a human-in-the-loop who makes the final decision, or other ethical concerns. For example, in e-commerce, although companies can assign different users to different visual interfaces of a website, they cannot compel users to sign up or subscribe. Instead, they can nudge or encourage users to sign up via promotions and offers. Encouragement designs are widely used in e-commerce to avoid withholding access to new features (Spotify Engineering 2023) and in social settings because “when evaluating a noncompulsory, entitlement program, researchers and practitioners cannot and should not force individuals to take up the program nor deny eligible individuals access to the program” (Heard et al. 2017).
Throughout, the decision-maker controls an encouragement or recommendation, not treatment uptake itself. Our analysis relies on a covariate-conditional exclusion restriction/assumption of no direct effect of encouragement on outcomes. Conditional on observed covariates, encouragement affects outcomes only through its effect on treatment take-up. This assumption encodes the “nudge” nature of encouragements, and it enables us to conceptually separate the effects of nudges on treatment take-up from the effects of treatments on downstream outcomes. To interpret that decomposition in terms of heterogeneous treatment effects, we additionally invoke treatment unconfoundedness; under truly randomized encouragement, alternative IV-style interpretations are also available. We therefore study policies that choose who to encourage so as to maximize policy value under the induced treatment behavior, optionally subject to budget and fairness constraints on induced treatment take-up or other groupwise impacts.
Optimal personalized policies often are optimizing in the space of encouragements, not treatments. Marginal encouragement effects are often enough for operational decision-making. For example, revenue management and pricing maximizes the expected profit, corresponding to the intention-to-treat effect. Operational goals to make better decisions differ from scientific goals to estimate the specific causal effect of receiving treatment, which gives rise to more challenging instrumental variables estimands. On the other hand, especially in the same consequential domains where it is impossible to compel treatment due to actions on and effects on people, decision-makers are interested separately in tracking who ultimately receives treatment (take-up) and who benefits from receiving treatment (treatment efficacy). For example, recent papers optimize for personalized allocation of job training services (Kallus 2023, Athey and Wager 2021, Bertsimas et al. 2019), leveraging data from large randomized experiments of job training programs (Imbens and Xu 2024). However, recent deployments of algorithmic job profiling in Europe have exposed key fairness concerns. A job profiling tool in Austria gave women and non-Austrian citizens lower predictions (hence lower priority for job training) and was ultimately scrapped by the data protection authority (Szigetvari ). The federal court overturned this decision, but the data had already been destroyed. The labor agency emphasized the predictive scores would only inform caseworker decisions, not dictate them. Would disparities in displayed predictive scores translate in practice also into discriminatory decisions for women, and worse labor market outcomes? Answering this requires separating how recommendations affect downstream human decisions from how those decisions affect outcomes.
Encouragement policy design is therefore inherently multi-objective: decision-makers may care about downstream outcomes, treatment take-up, encouragement burden, and disparities across groups. These tradeoffs may be represented either through penalized objectives or through constrained formulations.
The distinction between encouragements versus the treatments they nudge is ubiquitous in practice, arising whenever treatment assignment relies on intermediate human decisions. For example, price affects both demand and revenue. Demand fairness could be relevant (Cohen et al. 2022, Kallus and Zhou 2021a), while personalized pricing could expand access (Karlan and Zinman 2008). In healthcare, nonadherence is a major issue and can be significantly affected by social determinants of health (Miao et al. 2024). We later discuss a case study of targeted outreach for take-up of social services, so we discuss it here in more detail.
Example 1.1 (Takeup of social services and digital outreach)
In public programs such as Supplemental Nutrition Assistance Program (SNAP), many eligible individuals do not receive or retain benefits because of operational frictions in information, outreach, and benefits renewal (re-certification). These frictions are naturally modeled as barriers to take-up rather than as differences in treatment efficacy, which is fixed. Encouragements such as reminders, outreach, or navigation support are typically the only operational lever available to nonprofits and agencies to improve program enrollment among the likely-eligible. But positive average treatment effects hide groupwise heterogeneity where worse-off individuals may be less likely to act on nudges and enroll into beneficial treatment enrollment (Finkelstein and Notowidigdo 2019).
Our SNAP case study later returns to this setting and asks whether disparities under naive targeting are driven primarily by treatment benefit or by responsiveness to encouragement. We show that naive personalization can generate substantial access fairness gaps because responsiveness to encouragement is less aligned with treatment benefit for nonwhite recipients.
@@
Thus far, we have focused on modeling physically randomized encouragements such as those we have just described. However, our model can be extended to model the impacts of algorithmic advice on humans-in-the-loop. In this algorithmic advice setting, the covariate-conditional exclusion restriction is still plausible, but the “overlap” assumption is violated: Randomized (or as-if randomized) algorithmic encouragement or nudge recommendations are typically fixed functions of covariates. Crucially, our model’s decomposition of treatment take-up and effects on outcomes renders recent approaches based on robust optimization (Ben-Michael et al. 2021) less conservative.
Example 1.2 (Algorithmic advice)
@@
Doctors prescribe treatment on the basis of algorithmic recommendations (Lin et al. 2021), managers and workers combine their expertise to act on the basis of algorithmic decision support (Bastani et al. 2021), and in the social sector, caseworkers assign individuals to benefit programs on the basis of recommendations from risk scores that support triage (De-Arteaga et al. 2020, Green and Chen 2019, Yacoby et al. 2022).
A central managerial point is that disparities in encouragement settings can arise from different causal channels, including treatment efficacy and responsiveness to encouragement. Because these channels imply different operational responses, we advocate an audit-to-remedy workflow: first audit unconstrained encouragement policies by separately examining take-up and efficacy, then choose the appropriate intervention.
We study personalized decision problems in which the planner controls recommendations into treatment rather than treatment itself. Under a covariate-conditional no-direct-effect model of encouragement, policy value depends on two distinct objects: responsiveness to encouragement and treatment efficacy. This modeling distinction makes induced treatment take-up, rather than recommendation rates alone, the natural fairness target and yields tractable policy characterizations under budget and access constraints. In settings with deterministic algorithmic recommendations, the same model localizes overlap-robustness to the recommendation-response model rather than the downstream outcome model. In the supervised-release application, local changes to simple interpretable scorecards can more than halve current disparities in supervision at comparatively little value loss. Data-driven encouragement policies offer additional degrees of freedom for improving fairness upon the status quo.
@@ 2 Related Work
Fairness analysis of optimal policy learning and responses to algorithmic advice
Chohlas-Wood et al. (2021) puts forth a consequentialist framework for parity-aware policy design, emphasizing elicitation of preferences on efficiency and resource parity in a contextual bandit problem. In our work, the planner controls recommendations rather than treatment itself and disparities in induced take-up introduce another objective. The elicitation mechanisms of Chohlas-Wood et al. (2021) could similarly be used here to guide trade-offs on multiple objectives. Ge et al. (2025) studies how to robustly improve the algorithmic fairness of classifiers under downstream human decisions without any information on compliance. In contrast, our later extension to algorithmic advice introduces differential causal effects of receiving treatment, and throughout we assume observational data access to prior human-made decisions in response to classifiers. McLaughlin and Spiess (2022), Gillis et al. (2021) study the fairness of machine-assisted human decisions and develop a principal-agent model studying how algorithmic advice might shift beliefs in risk and preferences. In contrast, we are not micro-founded as to why disparities in compliance arise, which is an important direction of future work. Further, our assumptions are complementary and assume decision-makers are “set in their ways” responding to algorithmic risk labels: empirically supported by prior studies (Pruss 2023, Imai et al. 2020) in the criminal justice setting of our later case study. A rapidly growing literature in operations studies algorithmic advice, at times from a behavioral focus (but not specialized to causal effects). Grand-Clément and Pauphilet (2024) also shares a focus on leveraging adherence to optimize advice rather than decisions - however they adopt a dynamic model with updating.
Optimal encouragement designs/policy learning with constraints.
There is extensive literature on off-policy evaluation and learning, empirical welfare maximization, and optimal treatment regimes (Athey and Wager 2021, Zhao et al. 2012, Manski 2005, Kitagawa and Tetenov 2015). The most closely related work in terms of problem setup is the formulation of “optimal encouragement designs” in Qiu et al. (2021), but they focus on knapsack resource constraints. Sun et al. (2021) has studied uniform feasibility in constrained resource allocation, but without encouragement or fairness. Ben-Michael et al. (2021) studies robust extrapolation in policy learning from algorithmic recommendation, exactly implicitly optimizing in the space of encouragements, but not fairness concerns. By introducing separate consideration of treatment take-up and efficacy, we can substantially reduce the conservativity of robust formulations and limit extrapolation to behavioral response to recommendations rather than unknown treatment effects.
Fair pricing in operations: take-up (demand) vs. revenue
We view our perspective on fairness constraints as translating the normatively important role of demand fairness from fair pricing analysis in operations management to settings beyond revenue outcomes alone and our treatment access fairness framework. Cohen et al. (2022) study revenue and other implications of different potential fairness notions in pricing, including price, demand, consumer surplus, and no-purchase valuation under a linear demand model. Kallus and Zhou (2021a) study fairness in pricing, from a nonparametric, contextual perspective. Although fair pricing analysis often hinges specifically on the functional form of revenue and demand, its importance highlights the analogous importance of considering demand fairness in other decision-impact settings.
Algorithmic fairness in operations
Another rich line of work studies fairness in dynamic resource allocation (Manshadi et al. 2021, Salem et al. 2026, Freund et al. 2023, Banerjee et al. 2023). Our encouragements model points out that allocating resources need not ensure that individuals benefit: potentially due to concerning persistent access or take-up gaps. Garg et al. (2020) study how access gaps to test score information can change normative policy recommendations, Andrews et al. (2022) studies access gaps in resource allocation, and Garg (2025), Herd and Moynihan (2019) documents access gaps and administrative burdens in public-facing mechanisms.
3 Encouragement model and identification
We introduce the setup, key assumptions, and the resulting characterization of encouragement policies that underlies the analysis below, before discussing extensions via sensitivity analysis. The central modeling assumption is that recommendation affects outcomes only through treatment take-up. Under this encouragement model, policy value depends jointly on responsiveness to encouragement and treatment efficacy.
We work in the potential outcomes framework for causal inference (Rubin 1980), which posits that each individual is endowed with a vector of potential outcomes, one for each value of treatment. We define the following: